# Largest component in Boolean sublattices

### v1

@@ 2.1. The Lubell function
In the proof of Theorem 1.3, we relate the problem to another quantity. Let $\mathcal{F}\subseteq 2^{[n]}$. For $\sigma\in S_{n}$, call $T(\sigma)$ the number of times that $\sigma$ meets $\mathcal{F}$, that is the number of initial segments of $\sigma$ that are in $\mathcal{F}$. Then the Lubell function of $\mathcal{F}$ is
@@
$\overline{\Lambda}(n,\mathcal{F})=\frac{\sum_{i=1}^{n}s_{i}(\mathcal{F})}{s(\mathcal{F})}.$
@@ 2.3. Bounding $\xi^{\star}_{n}$
As we showed in Section 2.2, Lemma 2.5 allows us to use results about $\xi^{\star}_{n}$ to obtain results about $\operatorname{La}(n,t)$. We now relate $\xi^{\star}_{n}$ to a well-studied Turán-type problem. Recall that a proper edge-colouring of a graph $G=(V,E)$ is a mapping $c:E\to\mathbb{Z}^{+}$ such that any two incident edges $e_{1},e_{2}$ must have $c(e_{1})\neq c(e_{2})$. A cycle in $G$ of length $\ell\geq 3$ is called rainbow (with respect to $c$) if no two edges of the cycle receive the same colour. For $\ell\geq 3$, we denote by $\operatorname{ex}^{\star}(t,\mathcal{O}_{\ell})$, the maximum number of edges in a graph $G$ of order $t$ such that there exists a proper edge-colouring of $G$ with no copies of a rainbow cycle of length $\ell$. We denote by $\operatorname{ex}^{\star}(t,\mathcal{O}^{\star})$ the maximum number of edges in a graph $G$ of order $t$ such that there exists a proper edge-colouring of $G$ with no rainbow cycles.
@@
For all $n,t\geq 1$ we have $\xi^{\star}_{n}(t)\leq 2\operatorname{ex}^{\star}(t,\mathcal{O}^{\star})/t.$
@@
Assume for a contradiction that the inequality does not hold. This means there are two sets $\mathcal{A}\subseteq\binom{[n]}{k}$, $\mathcal{B}\subseteq\binom{[n]}{k+1}$ for some $0\leq k<n$ with $|\mathcal{A}|+|\mathcal{B}|=t$ and $\xi(\mathcal{A},\mathcal{B})>\operatorname{ex}^{\star}(t,\mathcal{O}^{\star})$. By definition of $\operatorname{ex}^{\star}(t,\mathcal{O}^{\star})$, this means that every proper edge-colouring of the graph $G_{\mathcal{A}\cup\mathcal{B}}$ contains a rainbow cycle. However, let us consider the following edge-colouring: given any edge $e$ of $G_{\mathcal{A}\cup\mathcal{B}}$, and supposing it corresponds to the pair $(A,A\cup\{i\})$ for some $i\in[n]$, then colour $e$ with colour $i$.
@@
Lemmas 2.7 and 2.5 gives us bounds on $\operatorname{La}(n,t)$ via the establishment of good bounds on $\operatorname{ex}^{\star}(t,\mathcal{O}^{\star})$. There has already been much work on this latter problem. This was initiated by Keevash, Mubayi, Sudakov and Verstraëte [16], who showed that $\operatorname{ex}^{\star}(t,\mathcal{O}^{\star})=O(t^{4/3})$, simply by bounding it with $\operatorname{ex}^{\star}(t,\mathcal{O}_{6})$. There have been numerous successive improvements [9, 12, 24, 17, 13]. The current state of the art is due to Alon, Bucić, Sauerman, Zakharov, Zamir [2].
@@
$\operatorname{ex}^{\star}(t,\mathcal{O}^{\star})=O(t\log t\log\log t)$.
@@ 5. Concluding remarks
Our main conjecture (1.1) proposes that the size guarantee $\exp(\Omega(n/\log n))$ above could be improved to $\exp(\Omega(n))$. (Consider the Stirling’s approximation computations displayed parenthetically after the conjecture.) This improvement would in fact follow if we could establish that $\operatorname{ex}^{\star}(t,\mathcal{O}^{\star})=O(t\log t)$, which is an open problem mentioned in [2], or, more specifically, if we could show $\xi^{\star}_{n}(t)=O(\log t)$, which is related to 2.6 and is an interesting problem in its own right. We contend that the special case $k=n-2$ of 1.1 alone is enticing, even only asymptotically. Moreover, 1.1 suggests that the size of a largest component will be of order $2^{n-o(n)}$ once we prescribe that $|\mathcal{F}|=\omega(\operatorname{La}(n))$. Thus, the broader picture sketched (if correct) compares interestingly against the intuition we might have borrowed from, say, the theory of random graphs or percolation.

### v2

@@ 2.1. The Lubell function
In the proof of Theorem 1.3, we relate the problem to another quantity. Let $\mathcal{F}\subseteq 2^{[n]}$. For $\sigma\in S_{n}$, call $T(\sigma)$ the number of times that $\sigma$ meets $\mathcal{F}$, that is $T(\sigma)=|\{F\in\mathcal{F}:\{\sigma(1),\sigma(2),\dots,\sigma(|F|)\}=F\}|$. Then the Lubell function of $\mathcal{F}$ is
@@
$\overline{\Lambda}(n,\mathcal{F})=\frac{\sum_{i=1}^{n}is_{i}(\mathcal{F})}{s(\mathcal{F})}.$
@@ 2.3. Bounding $\xi^{\star}_{n}$
As we showed in Section 2.2, Lemma 2.5 allows us to use results about $\xi^{\star}_{n}$ to obtain results about $\operatorname{La}(n,t)$. We now relate $\xi^{\star}_{n}$ to a well-studied Turán-type problem. Recall that a proper edge-colouring of a graph $G=(V,E)$ is a mapping $c:E\to\mathbb{Z}^{+}$ such that any two incident edges $e_{1},e_{2}$ must have $c(e_{1})\neq c(e_{2})$. A cycle in $G$ of length $\ell\geq 3$ is called rainbow (with respect to $c$) if no two edges of the cycle receive the same colour. For $\ell\geq 3$, we denote by $\operatorname{D}^{\star}(t,\mathcal{O}_{\ell})$, the largest average degree in a graph $G$ of order $t$ such that there exists a proper edge-colouring of $G$ with no copies of a rainbow cycle of length $\ell$. We denote by $\operatorname{D}^{\star}(t,\mathcal{O}^{\star})$ the largest average degree in a graph $G$ of order $t$ such that there exists a proper edge-colouring of $G$ with no rainbow cycles.
@@
For all $n,t\geq 1$ we have $\xi^{\star}_{n}(t)\leq\operatorname{D}^{\star}(t,\mathcal{O}^{\star}).$
@@
Assume for a contradiction that the inequality does not hold. This means there are two sets $\mathcal{A}\subseteq\binom{[n]}{k}$, $\mathcal{B}\subseteq\binom{[n]}{k+1}$ for some $0\leq k<n$ with $|\mathcal{A}|+|\mathcal{B}|=t$ and $2\xi(\mathcal{A},\mathcal{B})/t>\operatorname{D}^{\star}(t,\mathcal{O}^{\star})$. By definition of $\operatorname{D}^{\star}(t,\mathcal{O}^{\star})$, this means that every proper edge-colouring of the graph $G_{\mathcal{A}\cup\mathcal{B}}$ contains a rainbow cycle. However, let us consider the following edge-colouring: given any edge $e$ of $G_{\mathcal{A}\cup\mathcal{B}}$, and supposing it corresponds to the pair $(A,A\cup\{i\})$ for some $i\in[n]$, then colour $e$ with colour $i$.
@@
Lemmas 2.7 and 2.5 gives us bounds on $\operatorname{La}(n,t)$ via the establishment of good bounds on $\operatorname{D}^{\star}(t,\mathcal{O}^{\star})$. There has already been much work on this latter problem. This was initiated by Keevash, Mubayi, Sudakov and Verstraëte [16], who showed that $\operatorname{D}^{\star}(t,\mathcal{O}^{\star})=O(t^{1/3})$, simply by bounding it with $\operatorname{D}^{\star}(t,\mathcal{O}_{6})$. There have been numerous successive improvements [9, 12, 24, 17, 13]. The current state of the art is due to Alon, Bucić, Sauermann, Zakharov, Zamir [2].
@@
$\operatorname{D}^{\star}(t,\mathcal{O}^{\star})=O(\log t\log\log t)$.
@@ 5. Concluding remarks
Our main conjecture (1.1) proposes that the size guarantee $\exp(\Omega(n/\log n))$ above could be improved to $\exp(\Omega(n))$. (Consider the Stirling’s approximation computations displayed parenthetically after the conjecture.) This improvement would in fact follow if we could establish that $\operatorname{D}^{\star}(t,\mathcal{O}^{\star})=O(\log t)$, which is an open problem mentioned in [2], or, more specifically, if we could show $\xi^{\star}_{n}(t)=O(\log t)$, which is related to 2.6 and is an interesting problem in its own right. We contend that the special case $k=n-2$ of 1.1 alone is enticing, even only asymptotically. Moreover, 1.1 suggests that the size of a largest component will be of order $2^{n-o(n)}$ once we prescribe that $|\mathcal{F}|=\omega(\operatorname{La}(n))$. Thus, the broader picture sketched (if correct) compares interestingly against the intuition we might have borrowed from, say, the theory of random graphs or percolation.